OpenStudy (anonymous):

Write the solution in lowest terms. -1/3 + 3/4 =

OpenStudy (ghazi):

\[\frac{ -1 }{ 3 }+\frac{ 3 }{ 4 }=\frac{ -4+3}{ 4*3 }\]

OpenStudy (ghazi):

take the LCM of denominator